OpenStudy (anonymous):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? show your work plz

OpenStudy (anonymous):

T+B=56 B=8+2T Substitute B=8+2T into T+B=56 T+(8+2T)=56 3T+8=56 3T=56-8 T=48/3 T=16 Substitute T=16 into B=8+2T B=8+2(16) B=8+32 B=40

OpenStudy (anonymous):

Tony is 16 years old and Belinda is 40 years old.

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

You're welcome.